OpenStudy (anonymous):

Find 3 consecutive odd numbers where the product of the smaller two numbers is 46 less than the square of the largest number

OpenStudy (anonymous):

Let x = the lesser of the integers then x+2 and x+4 are the next two integers Solve\[x(x+2)=(x+4)^2-46\]\[x^2+2x=x^2+8x+16-46\]\[-6x=-30\]\[x=5\]Thus, the three consecutive odd integers are 5, 7, 9